# VoroUDF: Meshing Unsigned Distance Fields with Voronoi Optimization

### v1

@@ 3. Method
Unlike grid-based methods that subdivide space using a fixed structure (e.g., an adaptive octree) to locate cells intersecting $\chi$, lacking the flexibility to capture thin structures and often misses features at low resolutions (see Figure 4), our method begins by initializing $N$ movable seeds with gradients $\{\mathbf{x}_{i},\mathbf{n}_{i}\}_{i=1}^{N}$. These seeds are sampled uniformly within the bounding volume and projected onto the implicit iso-surface $\chi$ using their corresponding gradients. We then partition the domain into regions via a Voronoi decomposition induced by the seed set $\{\Omega_{i}\}_{i=1}^{N}$.
@@ 3.1. $L_{1}$ Tangent Energy
where $A$ is a $3\times 3$ matrix and $b$ is a $3\times 1$ vector (3 is the UDF dimension).
@@ 3.2. Feature-aware Repulsive Energy
However, allowing seeds to move freely within the space without considering geometric features can lead to undesirable effects. For example, seeds may drift completely away from the iso-surface $\chi$ during optimization. Inspired by Wang et al. (2025), we introduce a feature-aware gradient projection strategy to mitigate this issue using the geometric information extracted from the $L_{1}$ tangent energy defined in Sec. 3.1. More specifically, We examine the null space of $A$ in Equation 4 via singular value decomposition (SVD), writing $A=U\,\Sigma\,V^{T}$. If $\text{rank}(A)=r$, then the last $(\dim-r)$ columns of $V$ (denoted $V_{0}$) form an orthonormal basis for the null space of $A$. Our strategy is to project the force $\mathbf{g}^{i}=\sum_{j}\nabla E_{rep}^{ij}$ onto this null space:
@@
3.3. Mesh Reconstruction via the Dual of a Geodesic Voronoi Diagram
@@ B.3. More Comparisons on garment models
We further compare our method with seven state-of-the-art methods on (first) $100$ garment shapes from DeepFashion3D (Zhu et al., 2020) dataset. The results from seven state-of-the-art methods are generated at higher resolution ($2^{7}$), shown in Table 6, Figure 18 and Figure 19.

### v2

@@ 3. Method
Unlike grid-based methods that subdivide space using a fixed structure (e.g., an adaptive octree) to locate cells intersecting $\chi$, lacking the flexibility to capture thin structures and often misses features at low resolutions (see Figure 4), our method begins by initializing NN movable seeds with gradients $\{\mathbf{x}_{i},\mathbf{n}_{i}\}_{i=1}^{N}$. These seeds are sampled uniformly within the bounding volume and projected onto the implicit iso-surface $\chi$ using their corresponding gradients. We then partition the domain into regions via a Voronoi decomposition induced by the seed set $\{\Omega_{i}\}_{i=1}^{N}$.
@@ 3.1. $L_{1}$ Tangent Energy
where AA is a $3\times 3$ matrix and $b$ is a $3\times 1$ vector (3 is the UDF dimension).
@@ 3.2. Feature-aware Repulsive Energy
However, allowing seeds to move freely within the space without considering geometric features can lead to undesirable effects. For example, seeds may drift completely away from the iso-surface $\chi$ during optimization. Inspired by Wang et al. (2025), we introduce a feature-aware gradient projection strategy to mitigate this issue using the geometric information extracted from the $L_{1}$ tangent energy defined in Sec. 3.1. More specifically, We examine the null space of AA in Equation 4 via singular value decomposition (SVD), writing $A=U\,\Sigma\,V^{T}$. If $\text{rank}(A)=r$, then the last $(\dim-r)$ columns of $V$ (denoted $V_{0}$) form an orthonormal basis for the null space of AA. Our strategy is to project the force $\mathbf{g}^{i}=\sum_{j}\nabla E_{rep}^{ij}$ onto this null space:
@@
$3.$ Mesh Reconstruction via the Dual of a Geodesic Voronoi Diagram
@@ B.3. More Comparisons on garment models
We further compare our method with seven state-of-the-art methods on (first) 100100 garment shapes from DeepFashion3D (Zhu et al., 2020) dataset. The results from seven state-of-the-art methods are generated at higher resolution ($2^{7}$), shown in Table 6, Figure 18 and Figure 19.